A truck (6500 kg) is accelerating from 20.0 m/s to 30.0 m/s. calculate the work done?


1
Expert's answer
2021-04-06T18:25:29-0400

Answer


Average acceleration in time t

"a=\frac{30-20}{t}=\frac{10}{t}m/s^2"


Average distance travelled

"d=(30-20) t"


Work done

"W=6500\times \frac{10}{t}\times10t=650KJ"